what is the dimentions of a rectangle that has a perimeter 42 in. and an area of 104 sq in.
Answer by oberobic(2304)

P = perimeter
P = 2(L+W)
P = 42 in.
.
A = area
A = L*W
A = 104 sq. in.
.
L*W = 104
L = 104/W
.
substitute
.
2(104/W + W) = 42
.
divide both sides by 2
.
104/W + W = 21
.
multiply both sides by W to eliminate fraction
.
W(104/W + W) = 21W
.
104 + W^2 = 21W
.
subtract 21W from both sides
.
104 + W^2 -21W = 0
.
rearrange
.
W^2 -21W + 104 = 0
.
factor, if possible
.
(W -8)(W -13) = 0
.
Check with FOIL
W^2 -13W -8W + 104
OK
.
W = 8 or -13.
.
But a negative width is nonsense.
So, our only solution is W = 8.
.
Substitute W=8 to find L.
L*W = 104
L*8 = 104
L = 13
.
Check your work using the perimeter.
2(L+W) = 42 ?
2(8+13) = 2(21) = 42
Correct.
.
Answer: The dimensions of the rectangle are 8 in. and 13 in.
.
Done.